# Population Transformer: Learning Population-level Representations of Neural Activity

### v2

@@ 4 Experiment Setup
Table 2: Pretraining PopT is critical to downstream decoding performance (EEG data). We test on a seizure detection task (TUAB in Obeid & Picone (2016)) with 21 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against end-to-end pretrained EEG models in section 3 (values from the original works). Shown are the ROC-AUC mean and stdev across 5 random seeds. Best per section are bolded.
Figure 2: Compared to common aggregation approaches, pretrained PopT consistently yields better downstream decoding across tasks, data modalities, and temporal embedding types. NPopT = Non-pretrained PopT. (a) performance on four audio-linguistic iEEG tasks with 90 electrodes. Grey bars denote standard error across subjects. (b) performance on a seizure detection EEG task with 21 electrodes. Grey bars denote standard deviation across 5 random seeds.
@@ 5 Results
Decoding performance We find that using a pretrained PopT significantly benefits downstream decoding compared to baseline channel aggregation techniques across tasks, data modalities, and temporal encoding models (Tables 1, 2 and 2). The pretrained PopT scales well with increasing ensemble sizes (Figure 3a), a challenging task for the baseline aggregation approaches due to limited downstream task data and increasing input size. To test our method’s ability to handle multiple types of channel encodings, we applied our framework to 4 different channel encoders: (1) an iEEG-specific temporal encoder (BrainBERT (Wang et al., 2022)), (2) a general tokenization-based time-series encoder (TOTEM (Talukder et al., 2024)), (3) a pretrained general time-series encoder (Chronos (Ansari et al., 2024)), and a general convolution-based time-series encoder (TS2Vec (Yue et al., 2022)). We see significant improvements in performance with the pretrained PopT in all cases when comparing with baseline aggregation approaches (Figure 2).
@@
We also find that PopT can achieve competitive performance against pretrained end-to-end models, such as Brant (Zhang et al., 2024) for iEEG, and BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024) for EEG (Tables 1 and 2). For instance, PopT outperforms Brant (Zhang et al., 2024) in decoding iEEG data with our pretrained PopT + BrainBERT combination, likely due to PopT’s ability to leverage spatial relationships. Whereas Brant leaves the channel aggregation problem open. PopT is competitive with recent end-to-end trained EEG models (Yang et al., 2024; Jiang et al., 2024) at the EEG TUAB seizure detection task. This is impressive, since models such as LaBraM were specifically developed for this application, whereas PopT was trained on top of generic time-series embeddings. We find that PopT can offer an efficient and competitive alternative to large end-to-end models for these decoding tasks, due to the effectiveness of our pretraining task for learning spatial and functional relationships between channel input embeddings.
@@
To verify that the weights of the pretrained PopT capture neural processing well even without fine-tuning, we also train a linear-encoder on top of the frozen PopT [CLS] token and find the same trends (Figure 11. This point in particular is important in building confidence in the results of our interpretability studies (see Section 6), in which we use the frozen pretrained weights to analyze connectivity. Finally, for the remaining analyses described below, we use a PopT with BrainBERT inputs.
@@
Sample and compute efficiency Our PopT learns spatial relationships between channels, in a way that makes downstream supervised learning more data and compute efficient (see Figure 4 and Figure 5). Compared to the non-pretrained baseline models, fine-tuning the pretrained PopT can achieve the same decoding performance as other aggregation techniques with an order of magnitude fewer samples. The pretrained PopT surpasses the performance achieved by all other aggregation techniques by 500 samples out of the full dataset (roughly 5-10k examples depending on subject and task) Figure 4. The pretrained PopT also converges at a low number of steps, competitive with Linear and Deep NN compute efficiency. This greatly contrasts with the non-pretrained PopT and occasionally the Linear and Deep NN baselines, which may require 2k or more steps (Figure 5).
@@
Figure 5: Pretrained PopT is consistently compute efficient when fine-tuning. Number of steps required for each model to reach final performance during fine-tuning (dashed lines). The pretrained PopT consistently requires fewer than 750 steps (each step is an update on a batch size of 256) to converge, in contrast to the 2k steps required for the non-pretrained PopT. Linear and Deep NN aggregation can be similarly compute efficient, but occasionally require more training steps depending on task. Bands show standard error across subjects. Stars indicate fully trained performance.
Figure 6: Gains in decoding performance are available to new subjects. A minimal decrease in downstream decoding performance is found if the subject is held-out from pretraining (Held-out vs All). This is in stark contrast to the achievable downstream performance with a non-pretrained PopT (Non-pretrained PopT).
@@
Generalizability To test if our pretrained weights will be useful for subjects not seen during training, we conduct a hold-one-out analysis. We pretrain a model using all subjects except for one, and then fine-tune and evaluate on the model downstream. We find that missing a subject from pretraining does not significantly affect the downstream results (see Figure 6). This raises our confidence that the pretrained weights will be useful for unseen subjects and for researchers using new data.
@@
Scaling with number of pretraining subjects To investigate the effect of scaling pretraining data on our model, we pretrain additional versions of PopT using only 1, 2, or 3 subjects. We find a consistent improvement in downstream decoding when we increase the number of pretraining subjects available across all our downstream decoding tasks Figure 7. A significant improvement is found with just 1 pretraining subject already, potentially due to adaption to the temporal embeddings used. The decoding performance using all our pretraining data is significantly higher in most decoding tasks than with just 1 or 2 subjects in the pretraining data, suggesting the potential for our framework to continue scaling with more subjects.
Ablation of loss components and position information An ablation study confirms that both the ensemble-wise and channel-wise component of the pretraining objective contribute to the downstream performance (Table 3). Furthermore, including the 3D position information for each channel is critical for decoding. These findings also hold when the PopT is kept frozen during fine-tuning (see Appendix H: Frozen ablation). Additionally, we find that the discriminative nature of our loss is necessary for decoding. Attempting to add an L1 reconstruction term to our pretraining objective results in poorer performance, perhaps because the model learns to overfit on low-entropy features in the embedding. Our discriminative loss requires the model to understand the embeddings in terms of how they can be distinguished from one another, which leads the model to extract more informative representations.
@@
Figure 7: Pretraining with more subjects leads to better downstream performance. We pretrain PopT with different number of subjects (colors) and test on our decoding tasks (x-axis). Bars indicate mean and standard error of performance across channel ensembles 5-30 on a held out test subject. Pretraining with one subject gives a considerable benefit compared to no pretraining (red to yellow), but the addition of more subjects to pretraining consistently improves performance (yellow $\rightarrow$ green).
@@
PopT $\mathbf{0.69}\pm\mathbf{0.07}$ $\mathbf{0.84}\pm\mathbf{0.06}$ $\mathbf{0.86}\pm\mathbf{0.05}$ $\mathbf{0.89}\pm\mathbf{0.07}$
PopT w/o group-wise loss $0.66\pm 0.07$ $0.83\pm 0.06$ $0.84\pm 0.04$ $0.88\pm 0.08$
PopT w/o channel-wise loss $0.67\pm 0.06$ $0.81\pm 0.08$ $0.84\pm 0.06$ $0.87\pm 0.09$
PopT w/o position encoding $0.59\pm 0.07$ $0.67\pm 0.10$ $0.75\pm 0.08$ $0.79\pm 0.08$
PopT with reconstruction loss $0.60\pm 0.11$ $0.73\pm 0.11$ $0.81\pm 0.05$ $0.83\pm 0.09$
PopT with L1 reconstruction only $0.56\pm 0.04$ $0.65\pm 0.08$ $0.73\pm 0.10$ $0.74\pm 0.10$
Table 3: PopT ablation study. We individually ablate our losses and positional encodings during pretraining then decode on the resulting models. Shown are ROC-AUC mean and standard error across subjects evaluated at 90 electrodes. The best performing model across all decoding tasks uses all of our proposed components, showing that they are all necessary. Removing our positional encoding during pretraining and fine-tuning drops the performance the most, indicating that position encoding is highly important for achieving good decoding. Additionally, we attempt adding or only using a reconstruction component to the loss as a regularizing term, but find that this leads to poorer performance (last two rows).
Figure 8: Probing the pretrained model for inter-channel connectivity Traditionally, connectivity analysis between regions is done by computing the coherence between electrode activity (left). We propose an alternative analysis purely based on the contextual sensitivity learned during pretraining. Briefly, we select an electrode, mask out its activity, and then measure the degradation in the channel-wise objective function for the remaining electrodes. Plotting the values of this delta (right) recovers the main points of connectivity. Plots for all test subjects can be seen in Appendix F: Connectivity.
@@ 6 Interpreting Learned Weights
Finally, we conduct two interpretability studies of the Population Transformer’s learned weights.
Connectivity Traditional neuroscience analyses typically use cross-correlation as a measure of region connectivity (Wang et al., 2021). Our PopT allows for an alternative method of determining connectivity, based on the degree to which channels are sensitive to each other’s context. In this method, each channel is masked in turn, and then model performance on the pretraining channel-wise objective for the remaining unmasked channels is measured. We use the degradation in performance as a measure of connectivity. We can construct plots as in Figure 8, that recapitulate the strongest connectivity of the cross-correlation maps. Note that while some approaches for modelling brain activity explicitly build this into their architecture (Cai et al., 2023), we recover these connections purely as a result of our self-supervised learning. Additional method details available in Appendix E.
Candidate functional brain regions from attention weights After fine-tuning our weights on a decoding task, we can examine the attention weights of the [CLS] output for candidate functional brain regions. We obtain a normalized Scaled Attention Weight metric across all subjects to analyze candidate functional brain regions across sparsely sampled subject datasets Figure 9. The Scaled Attention Weight is computed from raw attention weights at the [CLS] token passed through the attention rollout algorithm (Abnar & Zuidema, 2020). The resulting weights from each channel are then grouped by brain region according to the Destrieux atlas (Destrieux et al., 2010). A full description of the method is available in Appendix E.
The resulting weights reveal expected functional brain regions related to the tasks decoded Figure 9, with low-level auditory tasks highlighting primary auditory cortex and higher-level language distinction tasks highlighting language-specific areas. Given the massive pretraining PopT undergoes, these scaled attention weights provide a valuable new tool for discovering candidate functional regions.
Figure 9: Attention weights from a fine-tuned PopT identify candidate functional brain regions Candidate functional maps can be read from attention weights of a PopT fine-tuned on our decoding tasks. For the Volume and Pitch tasks, note the weight placed on the primary auditory cortex (black arrows), but not in Wernicke’s area. For the Speech vs Non-speech and Sentence onset tasks, note the weight placed on regions near Wernicke’s area (black arrows). Center brain figure highlight regions related to auditory-linguistic processing (aph, 2017)).
@@ 7 Discussion
We presented a self-supervised scheme for learning effective joint representations of neural activity from temporal embeddings. Our approach improves decoding and reduces the samples required to learn downstream tasks, which is especially critical for neural data modalities given patient constraints. A key aspect of our approach is the fact that we focus on spatial aggregation of existing channel embeddings, rather than training a large end-to-end model. By decoupling temporal and spatial feature extraction, we are able to leverage existing temporal embeddings to learn spatiotemporal representations efficiently and with a smaller number of parameters. This makes our model available for use in low compute-resource settings. Furthermore, this separation of considerations opens up the possibility for future independent improvement in temporal modeling, whether that be from a domain specific model or a more general time-series encoder. The generality of this approach allowed us to train on two very different neural modalities: scalp EEG and invasive iEEG. Our success in these domains suggest that this approach could even be extended to settings outside of neuroscience that also contend with sparsely and variably distributed time-series data channels, as is often the case with geophysical or climate data.
Limitations and Future Work We propose an effective strategy for aggregating signals with meaningful spatial coordinates, but it remains to be seen how we can generalize this further to aggregations without such coordinates. Individual brains are also highly variable, so improving this positional encoding scheme could provide an even higher resolution latent space of each individual’s brain. Future work could experiment with automatic functional identification for each channel, such as that explored in neural spiking data (Azabou et al., 2024), but it is currently unclear how to do so with neural recordings that have lower SNR.
@@ 8 Conclusion
We introduced a pretraining method for learning representations of arbitrary ensembles of intracranial electrodes. We showed that our pretraining produced considerable improvements in downstream decoding and efficiency, that would not have been possible without the knowledge of spatial relationships learned during the self-supervised pretraining stage. These benefits were found across data modalities, decoding tasks, and temporal encoders used, speaking to the generality of our approach. We further showed that this scheme produces interpretable weights from which connectivity maps and candidate functional brain regions can be read. Finally, we release the pretrained weights for our PopT with BrainBERT inputs as well as our code for plug-and-play pretraining with any temporal embedding.
@@ 9 Acknowledgements
We would like to thank our funding sources: Caltech Chen Institute, Caltech Carver Mead New Adventures Fund. This work was also supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the NSF awards 2124052 and 2123818, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Artificial Social Intelligence for Successful Teams (ASIST) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0014, NIH grant R01EY026025, and the Office of Naval Research under award number N00014-20-1-2589 and award number N00014-20-1-2643. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We would also like to thank Xuefei Wang and Amy Li for helpful comments and discussion.

### v4

@@ 4 Experiment Setup
Table 2: Pretraining PopT is critical to downstream decoding performance (EEG data). We test on an abnormal EEG detection task (see Section 4) with 21 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against end-to-end pretrained EEG models in section 3 (values from the original works). Shown are mean and standard deviation across 5 random seeds. Best per section are bolded. Asterisks indicate that the bolded model is significantly better than the second-place model ($p<0.05$, Wilcoxon rank-sum).
Figure 2: Compared to common aggregation approaches, pretrained PopT consistently yields better downstream decoding across tasks, data modalities, and temporal embedding types. NPopT = Non-pretrained PopT. (a) performance on four audio-linguistic iEEG tasks with 90 electrodes. Grey bars denote standard error across subjects. (b) performance on an abnormal detection EEG task with 21 electrodes. Grey bars denote standard deviation across 5 random seeds.
@@ 5 Results
Decoding performance We find that using a pretrained PopT significantly benefits downstream decoding compared to baseline channel aggregation techniques across tasks, data modalities, and temporal encoding models (Tables 1 and 2 and Figure 2). To test our method’s ability to handle multiple types of channel encodings, we applied our framework to 4 different channel encoders: (1) an iEEG-specific temporal encoder: BrainBERT (Wang et al., 2022), (2) a general tokenization-based time-series encoder: TOTEM (Talukder et al., 2024), (3) a pretrained general time-series encoder: Chronos (Ansari et al., 2024), and a general convolution-based time-series encoder: TS2Vec (Yue et al., 2022). We see significant improvements in performance with the pretrained PopT in all cases when comparing with baseline aggregation approaches (Figure 2). Additionally, the pretrained PopT scales well with increasing ensemble sizes (Figure 3), a challenging task for the baseline aggregation approaches due to limited downstream task data and increasing input size.
@@
We also find that PopT can achieve competitive performance against pretrained end-to-end models, such as Brant (Zhang et al., 2024) for iEEG, and BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024) for EEG (Tables 1 and 2). For instance, PopT outperforms Brant (Zhang et al., 2024) in decoding iEEG data with our pretrained PopT + BrainBERT combination, likely due to PopT’s ability to leverage spatial relationships. Whereas Brant leaves the channel aggregation problem open. PopT is competitive with recent end-to-end trained EEG models (Yang et al., 2024; Jiang et al., 2024) on the EEG TUAB abnormal detection task. This is impressive, since models such as LaBraM were specifically developed for this application, whereas PopT was trained on top of generic time-series embeddings. We find that PopT can offer an efficient and competitive alternative to large end-to-end models for these decoding tasks, due to the effectiveness of our pretraining task for learning spatial and functional relationships between channel input embeddings.
@@
To verify that the weights of the pretrained PopT capture neural processing well even without fine-tuning, we also train a linear-encoder on top of the frozen PopT [CLS] token and find the same trends (Figure 17). This point in particular is important in building confidence in the results of our interpretability studies (Section 6), in which we use the frozen pretrained weights to analyze connectivity. For the remaining analyses described below, we use a PopT with BrainBERT inputs.
@@
Sample and compute efficiency Our PopT learns spatial relationships between channels, in a way that makes downstream supervised learning more data and compute efficient (Figure 4 and Figure 5). Compared to the non-pretrained baseline models, fine-tuning the pretrained PopT can achieve the same decoding performance as other aggregation techniques with an order of magnitude fewer samples. The pretrained PopT surpasses the performance achieved by all other aggregation techniques by 500 samples out of the full dataset (roughly 5-10k examples depending on subject and task) (Figure 4). The pretrained PopT also converges at a low number of steps. This greatly contrasts with the non-pretrained PopT. The Linear and Deep NN baselines can be similarly compute efficient, but occasionally may require 2k or more steps (Figure 5), as in the case of Speech vs. Non-speech.
@@
Figure 5: Pretrained PopT is consistently compute efficient when fine-tuning. Number of steps required for each model to reach final performance during fine-tuning (dashed lines). The pretrained PopT consistently requires fewer than 750 steps (each step is an update on a batch size of 256) to converge. Bands show standard error across subjects. Stars indicate fully trained performance.
@@
Generalizability To test if our pretrained weights will be useful for subjects not seen during training, we conduct a hold-one-out analysis. We pretrain a model using all subjects except for one, and then fine-tune and evaluate on the held-out subject. We find that missing a subject from pretraining does not significantly affect the downstream results (Figure 6). This raises our confidence that the pretrained weights will be useful for unseen subjects and for researchers using new data.
Figure 6: Gains in decoding performance are available to new subjects. A minimal decrease in downstream decoding performance is found if the subject is held-out from pretraining (Held-out vs All). Results are cross-validated across all test subjects. For BrainBERT, we report performance on the channel with the best linear-decodability. Markers show mean and standard error across subjects.
@@
Scaling with amount of pretraining data To investigate the effect of scaling pretraining data on our model, we pretrain versions of PopT using only 2%, 5%, 10%, 25%, and 50% of our data. Evaluation is performed on all test-subjects. We find a general improvement in downstream decoding when we increase the amount of pretraining data available across all our downstream decoding tasks (Figure 7), suggesting decoding improvements as we double our pretraining dataset. We expect that the framework could benefit from more diverse data, as a slight plataeuing effect is seen with the current pretraining dataset.
@@
Figure 7: Pretraining with more data leads to better downstream performance. We pretrain PopT with different percentages of our full pretraining dataset (colors) and test on our decoding tasks (x-axis). Markers show mean and standard error across test subjects.
Ablation of loss components and position information An ablation study confirms that both ensemble-wise and channel-wise losses contribute to downstream performance (Table 3). Furthermore, including the 3D position information for each channel and using discriminative losses are critical. Removing our positional encoding during pretraining and fine-tuning drops the performance significantly. Attempting to only use an L1 reconstruction term for our pretraining objective results in poorer performance. Our discriminative loss requires the model to understand the embeddings in terms of how they can be distinguished from one another, which leads the model to extract representations that are more beneficial for decoding.
@@
PopT $\mathbf{0.74}\pm\mathbf{0.03}$ $\mathbf{0.87}\pm\mathbf{0.03}$ $\mathbf{0.90}\pm\mathbf{0.01}$ $\mathbf{0.93}\pm\mathbf{0.02}$
PopT w/o channel-wise loss $0.71\pm 0.03$ $0.86\pm 0.03$ $0.88\pm 0.02$ $0.92\pm 0.02$
PopT w/o ensemble-wise loss $0.70\pm 0.03$ $0.87\pm 0.02$ $0.87\pm 0.01$ $0.92\pm 0.02$
PopT w/o position encoding $0.62\pm 0.07^{\vee}$ $0.76\pm 0.07^{\vee}$ $0.81\pm 0.09^{\vee}$ $0.83\pm 0.10^{\vee}$
PopT reconstruction loss only $0.56\pm 0.04^{\vee}$ $0.65\pm 0.08^{\vee}$ $0.73\pm 0.10^{\vee}$ $0.74\pm 0.10^{\vee}$
Table 3: PopT ablation study. We individually ablate our losses and positional encodings during pretraining then decode on the resulting models. Shown are ROC-AUC mean and standard error across subjects evaluated at 90 electrodes. The best performing model across all decoding tasks uses all of our proposed components. Here, $\vee$ denotes ablations which are significantly worse than the full model ($p<0.05$, Dunnett’s test).
@@ 6 Interpreting Learned Weights
Connectivity Traditional neuroscience analyses typically use cross-correlation as a measure of region connectivity (Wang et al., 2021). Our PopT allows for an alternative method of determining connectivity, based on the degree to which channels are sensitive to each other’s context. In this method, each channel is masked in turn, and then model performance on the pretraining channel-wise objective for the remaining unmasked channels is measured. We use the degradation in performance as a measure of connectivity. We can construct plots (Figure 8) that recapitulate the strongest connectivity of the cross-correlation maps. Note that while some approaches for modelling brain activity explicitly build this into their architecture (Cai et al., 2023), we recover these connections purely as a result of our self-supervised learning. Additional method details available in Appendix G.
Figure 8: Probing the pretrained model for inter-channel connectivity. Traditionally, connectivity analysis between regions is done by computing the coherence between electrode activity (left). We propose an alternative analysis purely based on the contextual sensitivity learned during pretraining. Briefly, we select an electrode, mask out its activity, and then measure the degradation in the channel-wise objective function for the remaining electrodes. Plotting the values of this delta (right) recovers the main points of connectivity. Plots for all test subjects can be seen in Appendix I.
Candidate functional brain regions from attention weights After fine-tuning our weights on a decoding task, we can examine the attention weights of the [CLS] output for candidate functional brain regions. We obtain a normalized Scaled Attention Weight metric across all subjects to analyze candidate functional brain regions across sparsely sampled subject datasets (Figure 9). The Scaled Attention Weight is computed from raw attention weights at the [CLS] token passed through the attention rollout algorithm (Abnar & Zuidema, 2020). The resulting weights from each channel are then grouped by brain region according to the Desikan-Killiany-Tourville (DKT) region atlas (Klein & Tourville, 2012). A full description of the method is available in Appendix G.
The resulting weights reveal expected functional brain regions related to the tasks decoded (Figure 9), with low-level auditory tasks highlighting primary auditory cortex and higher-level language distinction tasks highlighting language-specific areas. Given the training PopT undergoes, these weights provide a technique for discovering candidate regions (see Appendix H for quantitative comparison).
Figure 9: Attention weights from fine-tuned PopT identify candidate functional brain regions Candidate functional maps can be read from attention weights of a PopT fine-tuned on our decoding tasks. For the Speech vs Non-speech and Sentence onset tasks, note the weight placed on regions near Wernicke’s area (black arrows). In all tasks, the auditory cortex is attended to. Center brain figure highlight regions related to auditory-linguistic processing; figure credit: (aph, 2017)).
@@ 7 Discussion
We presented a self-supervised scheme for learning effective joint representations of neural activity from temporal embeddings. Our approach improves decoding and reduces the samples required to learn downstream tasks, which is especially critical for neural data modalities given subject constraints. A key aspect of our approach is the fact that we focus on spatial aggregation of existing channel embeddings, rather than training a large end-to-end model. By decoupling temporal and spatial feature extraction, we are able to leverage existing temporal embeddings to learn spatiotemporal representations efficiently and with a smaller number of parameters. This makes our model available for use in low compute-resource settings. Furthermore, this separation of considerations opens up the possibility for future independent improvement in temporal modeling, whether that be from a domain specific model or a more general time-series encoder. The generality of this approach allowed us to train on two very different neural modalities: scalp EEG and invasive iEEG. Our success in these domains suggest that this approach could even be extended to settings outside of neuroscience that also contend with sparsely and variably distributed time-series data channels, as is often the case with geophysical or climate data.
Limitations and Future Work We proposed a strategy for aggregating signals, provided that meaningful spatial coordinates are available, but it remains to be seen how to extend this approach to settings without such coordinates. Electrode layouts are highly variable, so it is important that some notion of positional encoding be given. Future work could experiment with automatic functional identification for each channel, such as that explored in neural spiking data (Azabou et al., 2024), but it is currently unclear how to do so with neural recordings that have lower SNR.
@@ 8 Conclusion
We introduced a pretraining method for learning representations of arbitrary ensembles of intracranial electrodes. We showed that our pretraining produced considerable improvements in downstream decoding and efficiency, that would not have been possible without the knowledge of spatial relationships learned during the self-supervised pretraining stage. These benefits were found across data modalities, decoding tasks, and temporal encoders used, speaking to the generality of our approach. We further showed that this scheme produces interpretable weights from which connectivity maps and candidate functional brain regions can be read. Finally, we release the pretrained weights for our PopT with BrainBERT inputs as well as our code for pretraining with any temporal embedding.
@@ 9 Acknowledgements
We would like to thank Xuefei (Julie) Wang and Kejun (Amy) Li for helpful comments and discussion.
We would like to thank our funding sources: Caltech Chen Institute, Caltech Carver Mead New Adventures Fund, Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the NSF award 2124052, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Mathematics for the DIscovery of ALgorithms and Architectures (DIAL) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0399, and the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.